# Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenario

### v1

@@ III.1 Walking dark QCD
III.2 QCD with ALP
One may also consider a new particle ($\Phi$) coupled to $M$, like $\Phi^{2}M^{2}$, and allow $\Phi$ to develop the vacuum expectation value and generate an extra soft-scale breaking mass term, to cancel the dynamical chiral symmetry breaking at low-enough $T$. Suppose that this vacuum expectation value becomes nonzero in the chiral symmetric phase at high $T$, and almost vanishes at the vacuum $T=0$. This setup ensures that the extra-soft scale breaking effect is activated to trigger the cosmological CW-type phase transition at $T=T_{n}$, and below $T_{n}$, it freezes out to make the QCD-chiral broken phase normal, so that the present-day QCD hadron physics is intact. The decoupling of the extra-soft scale breaking can be realized, say, if the vacuum expectation value of $\Phi$ scales with $\mu_{5}$, the chiral chemical potential, so the $\Phi$ field acts like an ALP coupled to the axial charge, as has been discussed in Zhang:2025kbu . This decoupling is then encoded with the freezing out of the QCD sphaleron, which would happen in the epoch of the confinement-deconfinement phase transition, which takes place seemingly simultaneously with the chiral phase transition. As long as the QCD sphaleron is active in hot QCD plasma, the chiral chemical potential $\mu_{5}$ along with a local CP-odd domain is created Manton:1983nd ; Klinkhamer:1984di , so that the QCD vacuum characterized by the strong CP phase $\theta$ and its fluctuation (in the spatial-homogeneous direction) gets significantly sizable Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe within the QCD time scale McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx : the sphaleron transition rate is not suppressed by the thermal effect in contrast to the instanton’s one Moore:1997im ; Moore:1999fs ; Bodeker:1999gx . The time fluctuation $\partial_{t}\theta(t)$, to be referred to as the chiral chemical potential, $\mu_{5}$ Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe ; Andrianov:2012hq ; Andrianov:2012dj , will be significant as well when the non-conservation law of the $U(1)$ axial symmetry is addressed.
We assume $\Phi$ to be a pseudoscalar which couples to up and down quarks via the Yukawa coupling term,
$y\,\bar{q}i\gamma_{5}\Phi q\,.$ (14)
Then, the $\Phi$ coupling to the chiral chemical potential $\mu_{5}$ would also be induced through the axial covariance:
$\frac{1}{2}|D_{\mu}\Phi|^{2}\,,\qquad{\rm with}\qquad D_{\mu}\Phi=(\partial_{\mu}-i\mu_{5}\delta^{0}_{\mu})\Phi\,.$ (15)
Hence $\Phi$ gets the effective potential along with $\mu_{5}$
$V_{\mu_{5}}(S)=-\frac{1}{2}\mu_{5}^{2}\Phi^{2}\,.$ (16)
This is definitely a negative mass-squared term ensured by the gauge covariance for the axialvector field $A_{\mu}=\mu_{5}\delta_{\mu}^{0}$, which corresponds to the repulsive force from the vector interaction against the scalar probes charged under the corresponding gauge. Thus, combined with the quartic $\lambda\Phi^{4}$ term, $\Phi$ can get nonzero vacuum expectation value, $v_{\Phi}$.
The Yukawa coupling term in Eq.(14), when put in the current MFA-NJL framework, modifies the thermodynamic potential. It makes a shift of the pseudoscalar mean field $P$ terms by $(y\,\Phi)$. Keeping $\theta=0$, we then find that up to thermal correction parts, the vacuum contribution in Eq.(10) looks like modified as
$\Omega$ $\approx\frac{\widetilde{M^{2}}}{2(g_{s}+g_{d})}+\frac{g_{d}\widetilde{P}^{2}}{g_{s}^{2}-g_{d}^{2}}-\frac{N_{f}N_{c}}{4\pi^{2}}\left[\widetilde{M}^{2}\Lambda^{2}-\frac{(\widetilde{M}^{2})^{2}}{4}\right]\ln\frac{\Lambda^{2}}{\widetilde{M}^{2}}\,,$ (17)
where
$\widetilde{M}^{2}$ $\equiv M^{2}+\widetilde{P}^{2}=4(g_{s}+g_{d})^{2}S^{2}+\widetilde{P}^{2}\,,$
$\widetilde{P}^{2}$ $\equiv[y\Phi-2(g_{s}-g_{d})P]^{2}\,.$ (18)
In Eq.(17) we have take the chiral limit $m\to 0$ for simplicity. The $\mu_{5}$-induced mass term for $\Phi$ in Eq.(16) makes the total potential aligned to develop the vacuum expectation values for $\Phi$ and also $P$, both of which are proportional to $\mu_{5}$. Hence the total potential will be stabilized separately along $M$ and $\widetilde{P}$ directions. Since the Yukawa coupling in Eq.(14) is CP invariant, we assume $\langle P\rangle=0$ at $\theta=0$, as in ordinary QCD. Then, in the $M$ direction, the developed $\langle\Phi\rangle_{\mu_{5}}$, generates the extra positive mass-squared contribution to $M$,
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}\langle\widetilde{P}\rangle_{\mu_{5}}^{2}\,,$ (19)
up to the logarithmic factor of ${\cal O}(1)$. Thus, the total $M^{2}$ term goes like
$\Omega_{M^{2}}$ $\sim\left[\frac{1}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\Lambda^{2}+\frac{N_{f}N_{c}}{8\pi^{2}}\langle y\,\Phi\rangle_{\mu_{5}}^{2}\right]M^{2}+\cdots\,,$
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}\left[\langle y\,\Phi\rangle_{\mu_{5}}^{2}-(475\,{\rm MeV})^{2}\right]M^{2}+\cdots\,,$ (20)
where ellipses stand for thermal corrections terms at $T\gtrsim T_{c}$, and we have quoted values of the model parameters referenced in Eq.(22). As clearly seen from Eq.(20), the soft-scale breaking mass term can be canceled when $\langle y\,\Phi\rangle_{\mu_{5}}\sim 475$ MeV, so that the chiral phase transition almost goes like the CW-type supercooling. It is interesting to note also that this $\Phi$-ALP, within only or extending the present framework can be a candidate to solve the strong CP problem.
The ALP-Yukawa coupling to light quarks as in Eq.(14) has severely been constrained by the astrophysical observations and collider experiments, via the ALP decay to diphoton. See the public website at https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html. The diphoton coupling for the present $\Phi$-ALP is induced by mixing with the QCD-$\eta^{\prime}$ meson, like in the case of composite ALP models at around the QCD scale Wang:2024vlm . The mixing arises through the Yukawa coupling $y$ in Eq.(14), so the $\Phi-\gamma-\gamma$ coupling $g_{\Phi\gamma\gamma}$ can roughly be of ${\cal O}(y\cdot m_{\eta^{\prime}}^{2}/|m_{\Phi}^{2}-m_{\eta^{\prime}}^{2}|\cdot g_{\eta\gamma\gamma})$, where $g_{\eta\gamma\gamma}\sim 10^{-3}\,{\rm GeV}^{-1}$ (see, e.g., Wang:2024vlm ) based on a chiral effective model approach. Since $\mu_{5}={\cal O}(T_{c})={\cal O}(100-200\,{\rm MeV})$, the size of $y$ is desired to be $\lesssim{\cal O}(1)$ and cannot be extremely small, it is likely that the $\Phi$-ALP is phenomenologically as heavy as $\gtrsim 1$ GeV, otherwise it has already been ruled out. More detailed phenomenological and cosmological consequences would be worth investigating, to be pursued elsewhere.
IV Exploring CW-type chiral phase transition:: case with $\mu_{B}\neq 0$

### v2

@@ III.1 The case with $\mu_{B}=0$
Below we will propose a new scenario at $\mu_{B}\sim 0$, which realizes the approximate scale invariance only at around $T=T_{c}$ (Sec. IV). Before this proposal, in the next section we shall take a look at the case with nonzero $\mu_{B}$, where the realization of the supercooling chiral phase transition, as has been discussed in the literature Shao:2024dxt ; Shao:2024ygm , will be recapped in terms of the class of the soft-scale breaking and the Debye mass effect.
III.2 The case with $\mu_{B}\neq 0$
@@
IV ALP assisted CW-type chiral phase transition in QCD
In this section, we propose a new scenario to realize the CW-type supercooling-chiral phase transition in the QCD thermal history. This scenario is categorized into the case with $\mu_{B}=0$.
We start with considering a new scalar field ($\Phi$) coupled to $M$, like $\Phi^{2}M^{2}$ of the portal form, and allow $\Phi$ to develop the vacuum expectation value and generate an extra soft-scale breaking mass term, to cancel the dynamical chiral symmetry breaking at low-enough $T$. Suppose that this vacuum expectation value becomes nonzero in the chiral symmetric phase at high $T$, and almost vanishes at the vacuum with $T=0$. This setup ensures that the extra soft-scale breaking effect is activated to trigger the cosmological CW-type phase transition at $T=T_{c}$, and below $T_{n}$, it freezes out to make the QCD-chiral broken phase normal, so that the present-day QCD hadron physics is intact.
The decoupling of extra soft-scale breaking can be realized, say, if the vacuum expectation value of $\Phi$ scales with $\mu_{5}$, the chiral chemical potential, so the $\Phi$ field acts like an ALP coupled to the axial charge, as has been discussed in Zhang:2025kbu . This decoupling is then encoded with the freezing out of the QCD sphaleron, which would happen in the epoch of the confinement-deconfinement phase transition, which takes place seemingly simultaneously with the chiral phase transition. As long as the QCD sphaleron is active in hot QCD plasma, the chiral chemical potential $\mu_{5}$ along with a local CP-odd domain is created Manton:1983nd ; Klinkhamer:1984di , so that the QCD vacuum characterized by the strong CP phase $\theta$ and its fluctuation (in the spatial-homogeneous direction) gets significantly sizable Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe within the QCD time scale McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx : the sphaleron transition rate is not suppressed by the thermal effect in contrast to the instanton’s one Moore:1997im ; Moore:1999fs ; Bodeker:1999gx . The time fluctuation $\partial_{t}\theta(t)$, to be referred to as the chiral chemical potential, $\mu_{5}$ Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe ; Andrianov:2012hq ; Andrianov:2012dj , will be significant as well when the non-conservation law of the $U(1)$ axial symmetry is addressed.
We assume an axial scalar field $\Phi=\phi_{s}+i\phi_{p}$, which we call the $\Phi$-ALP field, to couple to up and down quarks via the Yukawa coupling term,
$y\,(\bar{q}_{L}q_{R}\Phi+\bar{q}_{R}q_{L}\Phi^{*})=y(\bar{q}q\phi_{s}+\bar{q}i\gamma_{5}q\phi_{p})\,$, (16)
with $y\ll 1$. Then, the $\Phi$ coupling to the chiral chemical potential $\mu_{5}$ would also be induced through the $U(1)$ axial covariance:
$|D_{\mu}\Phi|^{2}\,,\qquad{\rm with}\qquad D_{\mu}\Phi=(\partial_{\mu}+2i\mu_{5}\delta^{0}_{\mu})\Phi\,.$ (17)
Hence $\Phi$ gets the effective potential along with $\mu_{5}$
$V_{\mu_{5}}(S)=-(2\mu_{5})^{2}|\Phi|^{2}\,.$ (18)
This is definitely a negative mass-squared term ensured by the gauge covariance for the external axialvector field $A_{\mu}=\mu_{5}\delta_{\mu}^{0}$, which corresponds to the repulsive force from the vector interaction against the scalar probes charged under the corresponding gauge.
Thus, combined with the quartic $\lambda|\Phi|^{4}$ term, $\Phi$ can get nonzero vacuum expectation value, $v_{\Phi}=\frac{2\mu_{5}}{\sqrt{\lambda}}$. We assume that the $\Phi$ field is settled at the vacuum expectation value at a high enough $T$, even when $T\gtrsim T_{c}$, where the QCD sphaleron actively generates $\mu_{5}$. At $T\sim T_{c}$, the Yukawa coupling term in Eq.(16), when put in the current MFA-NJL framework, modifies the thermodynamic potential to deform the type of the chiral phase transition. It makes a shift of the scalar ($S$) and pseudoscalar ($P$) mean fields terms only in the loop correction parts, by $(y\,\phi_{s})$ and $(y\,\phi_{p})$, respectively. Without loss of generality, we take $\langle\phi_{s}\rangle=v_{\Phi}$ and $\langle\phi_{p}\rangle=0$. Keeping $\theta=0$, hence $P=0$, we then find that up to thermal correction parts, the vacuum contribution in Eq.(10) looks like modified as
$\Omega$ $\approx\frac{{\cal S}^{2}}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\left[\widetilde{S}^{2}\Lambda^{2}-\frac{(\widetilde{S}^{2})^{2}}{4}\ln\frac{\Lambda^{2}}{\widetilde{S}^{2}}\right]\,$, (19)
where
$\widetilde{S}^{2}$ $\equiv(y\,v_{\Phi}-{\cal S})^{2}\,$,
${\cal S}^{2}$ $\equiv 4(g_{s}+g_{d})^{2}S^{2}\,$, (20)
In Eq.(19) we have taken the chiral limit $m\to 0$ for simplicity. Since the Yukawa coupling in Eq.(16) is CP invariant, we assume $\langle P\rangle=0$ at $\theta=0$, as in ordinary QCD. We find the extra positive mass-squared contribution to $\cal S$},
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}(yv_{\Phi})^{2}=\frac{N_{f}N_{c}}{2\pi^{2}}\frac{y^{2}\mu_{5}^{2}}{\lambda}\,$, (21)
up to the logarithmic factor of ${\cal O}(1)$. Thus, the total ${\cal M}^{2}$ term goes like
$\Omega_{M^{2}}$ $\sim\left[\frac{1}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\Lambda^{2}+\frac{N_{f}N_{c}}{8\pi^{2}}(yv_{\Phi})^{2}\right]{\cal S}^{2}+\cdots\,$,
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}\left[(yv_{\Phi})^{2}-(475\,{\rm MeV})^{2}\right]{\cal S}^{2}+\cdots\,$, (22)
where ellipses stand for thermal corrections terms at $T\gtrsim T_{c}$, and we have quoted values of the model parameters referenced in Eq.(15). As clearly seen from Eq.(22), the soft-scale breaking mass term can be canceled when $(yv_{\Phi})\sim 475$ MeV, so that the chiral phase transition almost goes like the CW-type supercooling.
Another type of enhancement of the positive soft-scale breaking would be generated from the intrinsic $\mu_{5}$ corrections acting as the catalyzer of the chiral/axial breaking under a proper regularization scheme Farias:2016let . Including this assist, the size of $(y\,v_{\Phi})$ desired to realize the CW-type supercooling would be smaller. The detailed study to precisely achieve the ALP-assisted CW-type supercooling is noteworthy to be performed in another publication.
At sufficiently low temperatures much below the confinement scale of ${\cal O}(100\,{\rm MeV})$, including the expected nucleation at $T_{n}$, where the theory is already also in the chiral broken phase with $\langle{\cal S}\rangle\simeq f_{\pi}\simeq 92$ MeV (where the estimated value follows from input parameters in Eq.(15) with the estimated MF quark condensate $S\equiv\langle\bar{q}q\rangle\simeq(-330\,{\rm MeV})$). Then the $\mu_{5}$ terms will be gone because of decoupling of the QCD sphaleron, consequently, $(y\,v_{\Phi})$ also goes away. Still, however, the $\Phi$-ALP field gets negative-mass squared correction from the quark loop, enough to develop its vacuum expectation value in the confining and chiral broken phase. The thermodynamic potential in Eq.(19) indeed provides an effective potential for $\phi=(y\phi_{s})$ like
$\Omega(\phi=y\phi_{s})\Bigg|_{T\ll T_{c}}\sim-\frac{N_{c}N_{f}}{4\pi^{2}}\Lambda^{2}(f_{\pi}-\phi)^{2}+\frac{N_{f}N_{c}}{16\pi^{2}}(f_{\pi}-\phi)^{4}\ln\frac{\Lambda^{2}}{(f_{\pi}-\phi)^{2}}\,.$ (23)
This, together with the quartic coupling term $\lambda|\Phi|^{4}=(\lambda/y^{4})|\phi|^{4}$, yields the vacuum expectation value of $\Phi$, $(v_{\Phi})_{0}$,
$(v_{\Phi})_{0}\approx-\left(\frac{y^{4}}{\lambda}\frac{N_{c}N_{f}}{2\pi^{2}}f_{\pi}\Lambda^{2}\right)^{1/3}\,$, (24)
for $(v_{\Phi})_{0}\ll f_{\pi}$. Here we have assumed $y^{4}\ll\lambda$, so that the $\lambda$-quartic term highly dominates over the loop-induced quartic term. This $(v_{\Phi})_{0}$ is generated essentially by the explicit breaking, via the $y$-Yukawa coupling term in Eq.(16) with $\langle\bar{q}q\rangle\neq 0$.
Around the vacuum at $(v_{\Phi})_{0}$, the $\Phi$-ALP field thus fluctuates to develop the mass along its radial component $\chi$ in the polar decomposition form,
$\Phi=\chi\cdot e^{ia/(v_{\Phi})_{0}}\,.$ (25)
The $\chi$-ALP mass $m_{\chi}$ is then evaluated as
$m_{\chi}\approx\frac{\sqrt{\lambda}}{y}|(v_{\Phi})_{0}|$ $\approx\left(\frac{\lambda^{1/2}}{y}\frac{N_{c}N_{f}}{2\pi^{2}}f_{\pi}\Lambda^{2}\right)^{1/3}$
$\simeq 214\,{\rm MeV}\times\left(\frac{\lambda^{1/2}}{y}\right)^{1/3}\,$, (26)
for $N_{c}=3$ and $N_{f}=2$. Here we have ignored the quark-loop induced wavefunction renormalization correction. The Nambu-Goldstone boson mode $a$ in $\Phi$ also gets the mass and the axionlike potential due to the generation of the explicit-breaking tadpole induced by $\langle\bar{q}q\rangle\neq 0$ in Eq.(16):
$V(a)=y(-v_{\Phi})_{0}\langle-\bar{q}q\rangle\cos\frac{a}{(v_{\Phi})_{0}}\,.$ (27)
From this potential, the $a$-ALP mass reads
$m_{a}^{2}=\frac{y\langle-\bar{q}q\rangle}{(-v_{\Phi})_{0}}$ $\simeq\frac{yf_{\pi}}{2(g_{s}+g_{d})(v_{\Phi})_{0}}$
$\simeq\left[124\,{\rm MeV}\times\left(\frac{\lambda}{y}\right)^{1/6}\right]^{2}\,$, (28)
where we have used the MFA $S\equiv\langle\bar{q}q\rangle$, which equals to ${\cal S}/(2(g_{s}+g_{d}))$ following from Eq.(20). Note that $m_{a}<m_{\chi}$ for any of $\lambda$ and $y$, as long as $y\ll 1$ and $\lambda={\cal O}(1)$.
The ALP-Yukawa coupling to light quarks as in Eq.(16) has severely been constrained via the ALP decay to diphoton by the astrophysical, cosmological, and collider experimental observations. See the public website at https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html. The diphoton coupling for the $\Phi$-ALP field is generated by mixing with the QCD-$\sigma$ (for the $\chi$-ALP) and $\eta^{\prime}$ mesons (for the $a$-ALP), arising through the Yukawa coupling $y$ in Eq.(16). We may roughly evaluate the $\chi(a)-\gamma-\gamma$ coupling at the $\chi/a$ onshell, $g_{\chi(a)\gamma\gamma}$, as
$g_{\chi(a)\gamma\gamma}\approx\frac{y\,\Lambda^{2}}{\sqrt{(m_{\chi/a}^{2}-m_{f_{0}(500)/\eta^{\prime}}^{2})^{2}+m_{f_{0}(500)/\eta^{\prime}}^{2}\Gamma^{2}_{f_{0}(500)/\eta^{\prime}}}}\times g_{f_{0}(500)/\eta^{\prime}\gamma\gamma}\,$, (29)
where $\Gamma_{f_{0}(500)/\eta^{\prime}}$ denotes the total widths of $f_{0}(500)$ and $\eta^{\prime}$, respectively. In Eq.(29) the denominator form comes from the $f_{0}(500)\eta^{\prime}$ propagator in the Breit-Wigner form, evaluated at the transfer momentum squared $q^{2}=m_{\chi/a}^{2}$, and the model-dependent factor $\sim y\Lambda^{2}$ denotes the mass mixing strength between the ALPs and those mesons. In the present evaluation, the latter part has been derived from the MFA-NJL based on the effective potential form in Eq.(19), with $\cal S$} and $v_{\Phi}$ replaced by $\eta^{\prime}$ and $(v_{\Phi})_{0}\sin a/(v_{\Phi})_{0}$ following from the polar decomposition of $\Phi$ in Eq.(25) applied to the $y$-Yukawa term in Eq.(16). The wavefunction renormalization of $f_{0}(500)/\eta^{\prime}$ has also been taken into account, which in the NJL-MFA yields $\sim N_{c}N_{f}/(16\pi^{2})\ln\Lambda^{2}/f_{\pi}^{2}$ for the inverse propagator in the chiral broken phase with $(v_{\Phi})_{0}\ll f_{\pi}$. Although the present NJL model only includes two flavors, the essential cutoff $\Lambda$ dependence on the mixing strength and wavefunction renormalization factor should be generic up to a factor of ${\cal O}(1)$, for the flavor singlet mesons, such as $\eta_{0}$ for the two-flavor case and $\eta^{\prime}_{0}$ for the three-flavor case. Therefore, the thus estimated net model-dependent factor $\sim y\Lambda^{2}$ in Eq.(29) covers the sufficient enough size in order of magnitude.
The $f_{0}(500)$ coupling to diphoton has not yet been clearly determined in experiments ParticleDataGroup:2024cfk . Moreover, the $\chi$ resonance will be contaminated with the $f_{0}(500)$ production, e.g., in the $\pi\pi$ scattering process. This channel will thus be challenging to use the probe of the present ALP-assisted model. Therefore, at this point, we disregard discussing the phenomenological constraint and consequence of $\chi$.
The total width of $\eta^{\prime}$ is available from the Particle Data Group ParticleDataGroup:2024cfk ; at the central value, $\Gamma_{\eta^{\prime}}\simeq 0.188$ MeV at $m_{\eta^{\prime}}\simeq 958$ MeV. We neglect the small mixing between $\eta^{\prime}$ and $\eta$. Regarding the $\eta^{\prime}$ mixing-induced ALP-photon coupling, in that case, we may refer to the case of composite ALP models at around the QCD scale Wang:2024vlm , where $g_{\eta\gamma\gamma}\sim 10^{-3}\,{\rm GeV}^{-1}$, which corresponds to the soft-$\eta^{\prime}$} limit value based on a chiral effective model approach.
Figure 3: The predicted curve of $g_{a\gamma\gamma}$ as a function of $m_{a}$ with $\lambda=1$ fixed. The plot range has taken into account the lower bound on $m_{a}$, set when $y=\lambda$ to ensure the perturbatively small enough quark loop correction to the ALP sector. The spike structure is generated at the resonance point when $m_{a}=m_{\eta^{\prime}}$, based on the resonant form factor structure of $g_{a\gamma\gamma}$ in Eq.(30). All the domains plotted in this figure survive over the current experimental constraint summarized in the webcite https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html.
Thus, we focus on the $a-\gamma-\gamma$ coupling,
$g_{a\gamma\gamma}$ $\approx\frac{y\,\Lambda^{2}}{\sqrt{(m_{a}^{2}-m_{\eta^{\prime}}^{2})^{2}+m_{\eta^{\prime}}^{2}\Gamma^{2}_{\eta^{\prime}}}}\times g_{\eta^{\prime}\gamma\gamma}\,.$ (30)
Solving Eq.(28) with respect to $y$; $y=(124\,{\rm MeV}/m_{a})^{6}\lambda$, and putting this $y$ into Eq.(30), we numerically write $g_{a\gamma\gamma}$ as a function of $m_{a}$. This function draws the theoretical prediction curve on the plane $(m_{a},g_{a\gamma\gamma})$ available at the AxionLimits (https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html). The predicted curve on $(m_{a},g_{a\gamma\gamma})$ is plotted in Fig. 3, where we have taken $\lambda=1$. We find that when $m_{a}\gtrsim 32\,{\rm MeV}$, where $g_{a\gamma\gamma}\lesssim 10^{-12}\,{\rm GeV}^{-1}$, the curve can be consistent with the current experiments, astrophysical, and cosmological observations. However, this lower bound on $m_{a}$ gives too large $y$: $y\sim\lambda$ in magnitude, therefore, the actual lower bound is set when $y=\lambda$, i.e., $m_{a}\simeq 124$ MeV. Figure 3 has already taken into account this limit. As a benchmark, we have $g_{a\gamma\gamma}\sim 10^{-17}$ for $m_{a}\sim 1$ GeV, which corresponds to the case with $y\sim 10^{-6}$ with $\lambda=1$ fixed.
This heavy ALP is unlikely to be probed or excluded even by the upcoming prospected experiments, such as the Electron Ion Collider experiment Balkin:2023gya . However, the indirect evidence could be tested via the GW and PBH productions sourced from the CW-type supercooling chiral phase transition, as has been suspected above. Detailed studies on this kind of cosmological observation feasibility are to be pursued in another publication.